OpenStudy (anonymous):

3x+2y=10 2x+5y=3 using the addition method

OpenStudy (anonymous):

someone please help me

OpenStudy (radar):

Assuming the addition method and the elimination method are the same thing. You would add the two expressions so that one of the variables will drop out (become zero) and then solve for the remaining variable. In order for the variable to "drop out" the coefficient of each must be equal in value and opposite in sign (in order to cancel or become zero. Do you understand what needs to be accomplished?

OpenStudy (radar):

Firs choose which variable you wish to get rid of..............x or y???

OpenStudy (radar):

I hereby choose x.

OpenStudy (radar):

Multiply first equation by -2 (all terms) getting -6x-4y=-20 Multiply second equation by 3 getting: 6x +15y = 9 Now we add.

OpenStudy (radar):

-6x-4y=-20 6x +15y=9 ------------ 11y=-11 y = -1 Note the x disappeared as the coefficient of x became zero (0) Now it is very easy to solve for x by substituting -1 for y in one of the original equations.